OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 22 and the last term is 369,754? Answers to choose from. a)372,552 b)392,160 c)411,768 d)431,376

OpenStudy (anonymous):

Term n = ar^n-1 This gives you the value of the nth term Sub in n=6, a=22 and Tn = 359754

OpenStudy (anonymous):

Sum of nth term = (a(r^n -1))/(r-1) Sub in values to find sum

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hey could you please plug the numbers in and ill multiply im a little confused @Skaematik

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

tn = a r^n-1 a = 22 n=6 Tn = 359754

OpenStudy (anonymous):

sn = (a(r^n -1))/(r-1) a = 22 n=6 sn= ? r= find r from previous formula

OpenStudy (anonymous):

thats the problem im having is how to plug the numbers in @Skaematik

OpenStudy (anonymous):

trying plugging them

OpenStudy (anonymous):

i Have and the answer i got was 411,768 is that correct? @Skaematik

OpenStudy (anonymous):

You need to use the exact value of the r value. Do not round up or down. I got 420094

OpenStudy (anonymous):

ohh i see thankyou very much

OpenStudy (anonymous):

@sauravshakya can you help me

OpenStudy (anonymous):

@bloodrain

OpenStudy (anonymous):

The answer you got was not any of the ones listed @Skaematik

OpenStudy (anonymous):

431376

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[369754 = 22 * r ^{6-1}\]\[S _{n}=\frac{ a \left( r ^{n}-1 \right) }{ r-1}\]

OpenStudy (anonymous):

thank you